OpenStudy (anonymous):

Martha estimated there were 99 marbles in a jar for a contest. The actual number of marbles in the jar was 121. What was the percent error of Martha's estimation?

OpenStudy (anonymous):

22% 18.18% 81.82% 22.22%

OpenStudy (anonymous):

idk how to do this problem can you explain it more

OpenStudy (anonymous):

\[\frac{ 121-99 }{ 121 }\times 100=?\]

OpenStudy (anonymous):

ok here is what you do now you have to do this equation which should give you the answer -99 times 100